### v2

@@ 1 Introduction
The gyroscope is essential equipment for measuring angular velocity in a wide range of technologies, such as motion tracking Scapellato et al. (2005); Chen et al. (2020), vibration measurement Yu and Pan (2017); Halkon and Rothberg (2021), and wearable health monitoring Cao and Sepúlveda (2019); Petropoulos et al. (2020); Farooq and Zhang (2022). Recently, we designed a low-cost micro-electromechanical (MEMS) inertial measurement unit (IMU) device to be worn by preganant women during childbirth to explore the impact of hospital birth room configuration upon mobility for women with complex pregnancies (See Figure. 1). However, a key concern of measurement in such ecological settings is that the accuracy of low-cost gyroscopes are usually low. For instance, when calculating the attitude, the integration will lead to the accumulation of drift and scale factor error Zhang et al. (2019). In addition, owing to the poor repeatability and significant volatility, on every booting or under different environmental conditions, the scale factor and biases change Yang et al. (2018); Zhang et al. (2017). Therefore, it is necessary to calibrate the gyroscope on each initialization or when environmental conditions change. However, calibration processes are normally time-consuming or cost ineffective, and thus in field or clinical context, practitioners need frequent calibration of the gyroscope, which needs to be a simple and efficient process.
@@ 1.1 Preliminary Study
To improve the estimation accuracy, we tried to develop in-field calibration methods for both accelerometers and gyroscopes. Despite the extensive amount of literature investigating in-field calibration of triaxial accelerometers or magnetometers Won and Golnaraghi (2009); Ye et al. (2017); Gao et al. (2017); Sipos et al. (2011); Wu et al. (2021); few studies discuss the calibration of the gyroscope under non-laboratory environments. This gap exists because local gravity serve as direct calibration references for accelerometers. However, when it comes to gyroscopes, the Earth’s self-rotation speed is insufficient as a standalone reference for accurate calibration. In this study, we introduce a novel, easy-to-use method for gyroscope calibration that significantly improves angle estimation accuracy, as demonstrated in subsequent discussions.
@@ 1.2 Existing Gyroscope Calibration Methods
Existing research recognized the critical role played by the methods and accuracy of gyroscope calibration. In particular, the conventional calibration method uses a turntable to provide a standard rotation speed for the gyroscope Chatfield (1997). This method can provide high calibration accuracy but requires expensive and precise instruments and complex calibration procedures that preclude its use in consumer electronic devices or clinical settings. Several recent studies Hwangbo et al. (2013); Wu and Pei (2017); Olivares et al. (2009); Ren et al. (2015); Lu et al. (2022); Xia et al. (2022) proposed several calibration methods that do not need precision equipment. Specifically, in Hwangbo et al. (2013), a camera-aided calibration method was reported. The images provide the orientation and position information of the sensor to confirm its orientation, resulting in a high computational complexity. In a separate method Wu and Pei (2017), a homogeneous magnetic field was employed as the calibration reference. The natural geomagnetic field is very weak and is easily affected by the alternating electric field, again making it difficult to implement outside the laboratory. In Olivares et al. (2009); Ren et al. (2015), an accelerometer-aided gyroscope calibration method was presented. The accelerometer was first calibrated using the multi-position method. Then, the rotation speed of the sensor body is provided by the accelerometer. In this method, an extra triaxial accelerometer with the same coordinate system is needed, and the calibration error of the accelerometer affects the calibration accuracy of the gyroscope. Besides, the calibration process takes more than ten minutes. In Lu et al. (2022), a two-step method was proposed to estimate all parameters for a six-axis IMU. A turntable was employed to perform a 12-points rotation, and a Kalman filter was designed to compensate for the parameter error. However, from the above discussion, we can infer that these gyroscope calibration methods relying on external equipment are unsuitable for scenarios where the external calibration device is unavailable, such as in a field or clinical setting akin to a busy and chaotic hospital birthing room. Besides, an accelerometer may not necessarily be integrated into the device and cannot provide calibration reference for a gyroscope. This paper, for the first time, introduced an easy-to-get rotation angle that can be used as the calibration reference. In this case, the gyroscope calibration no longer requires a constant rotation speed or reference from an accelerometer, thereby getting rid of the shackles of high-precision external equipment.
@@ 2 Methodology
In this section, we first propose an autocalibration method that uses manual rotation angle as a calibration reference. Then, a four-observation calibration method is reported. In this study, we ignored the earth’s rotation as it is often submerged in the noise of the low-cost gyroscope.
@@ 2.2 Calibration principle
where the $k_{x},k_{y},k_{z}$ and $b_{x},b_{y},b_{z}$ represents the scale factors and biases, respectfully.
@@
$R_{t}=(cos\theta_{t})I+(sin\theta_{t})[U_{t}]_{x}+(1-cos\theta_{t})U_{t}U_{t}^{T},$ (4)
where $I$ is a 3-by-3 identity matrix, and $[U_{t}]_{x}$ is the cross product matrix. We consider the orientation of the gyroscope to start reading (i.e. $t=0$) as the initial frame (${}^{n}$), and the gyroscope as body frame (${}^{b}$). Then, the rotation matrix $C^{n}_{b,t}$ which transforms the body frame to initial frame can be calculated as follows:
@@
where $K=[k_{x},k_{y},k_{z}]^{T}$ contains the scale factors. Let
@@
where $\theta_{total}{t}$ is a known value, $\tilde{C}_{t}^{n}$ is a constant value calculated by (8), and $t$ is the time of collecting $N$ samples, and $\theta_{total}$ is the total rotation angle performed in each experiment.
For an embedded system, time $t$ can be measured precisely. Turning a precise angle in the experiment becomes the key to the success of the experiment.
@@
Gyroscope readings $M_{t}=[m_{x,t},m_{y,t},m_{z,t}],i\in[0,T]$ for the stationary stage ($S_{0}$) and three rotating stages ($S_{1},S_{2},S_{3}$) (See Fig. 5). $i=1$, $K^{0}=[1,1,1].$
@@ 3 Simulation
Based on the above assumptions, we formulate the following simulation process. 30 sets of parameters are randomly generated based on the assumptions above. For each set of parameters, 500 simulations were repeated. During each simulation, we created a 4-observation measurement according to the experiment scheme described in Section 2.3. For the purpose of showing that the proposed method is not sensitive to the speed variation, the rotation speed in the rotating stage was represented by a randomly generated Bézier curve Han et al. (2008). A typical run of one simulation is shown in Figure. 6. We also generated a testing set for each simulation. The gyroscope were assumed to be rotating constantly and have equal speed projection on each axis during testing process.
@@ 4.2 Experimental Setting
calibration method turntable method Chatfield (1997) -aided method Ren et al. (2015) -aided method Wang et al. (2021)
@@
calibration method turntable method Chatfield (1997) -aided method Ren et al. (2015) -aided method Wang et al. (2021)
@@ 4.3 Comparing with existing calibration methods
To show the efficiency and accuracy of our method, we compared our method with two state-of-the-art autocalibration method Ren et al. (2015); Wang et al. (2021) and the gold standard turntable method Chatfield (1997).
@@
The accelerometer-aided method. This method uses the feature that most IMUs include accelerometers, and uses accelerometers to provide a calibration reference for the calibration of the gyroscope. Following the experiment process described in Ren et al. (2015), we reproduced the experiment using our gyroscopes. In our experiment, the inclination angle between the device and the desktop is set to 0°, 30°, and 45 °. Different from the original experiment, we did not repeat the experiment multiple times and take average to improve the calibration accuracy, because the experiment time is already as long as ten minutes.
The servomotor-aided method. This method uses a servomotor to provide a reference for calibration and provides a fast and low-cost calibration method. Due to the use of external calibration equipment, this method is not a pure autocalibration method. We have calibrated our gyroscope with reference to the calibration method provided in Wang et al. (2021).

### v3

@@ 1 Introduction
The gyroscope is essential equipment for measuring angular velocity in a wide range of technologies, such as motion tracking (Scapellato et al., 2005; Chen et al., 2020), vibration measurement (Yu and Pan, 2017; Halkon and Rothberg, 2021), and wearable health monitoring (Cao and Sepúlveda, 2019; Petropoulos et al., 2020; Farooq and Zhang, 2022). Recently, we designed a low-cost micro-electromechanical (MEMS) inertial measurement unit (IMU) device to be worn by preganant women during childbirth to explore the impact of hospital birth room configuration upon mobility for women with complex pregnancies (See Figure. 1). However, a key concern of measurement in such ecological settings is that the accuracy of low-cost gyroscopes are usually low. For instance, when calculating the attitude, the integration will lead to the accumulation of drift and scale factor error (Zhang et al., 2019). In addition, owing to the poor repeatability and significant volatility, on every booting or under different environmental conditions, the scale factor and biases change (Yang et al., 2018; Zhang et al., 2017). Therefore, it is necessary to calibrate the gyroscope on each initialization or when environmental conditions change. However, calibration processes are normally time-consuming or cost ineffective, and thus in field or clinical context, practitioners need frequent calibration of the gyroscope, which needs to be a simple and efficient process.
@@ 1.1 Preliminary Study
To improve the estimation accuracy, we tried to develop in-field calibration methods for both accelerometers and gyroscopes. Despite the extensive amount of literature investigating in-field calibration of triaxial accelerometers or magnetometers (Won and Golnaraghi, 2009; Ye et al., 2017; Gao et al., 2017; Sipos et al., 2011; Wu et al., 2021); few studies discuss the calibration of the gyroscope under non-laboratory environments. This gap exists because local gravity serve as direct calibration references for accelerometers. However, when it comes to gyroscopes, the Earth’s self-rotation speed is insufficient as a standalone reference for accurate calibration. In this study, we introduce a novel, easy-to-use method for gyroscope calibration that significantly improves angle estimation accuracy, as demonstrated in subsequent discussions.
@@ 1.2 Existing Gyroscope Calibration Methods
Existing research recognized the critical role played by the methods and accuracy of gyroscope calibration. In particular, the conventional calibration method uses a turntable to provide a standard rotation speed for the gyroscope (Chatfield, 1997). This method can provide high calibration accuracy but requires expensive and precise instruments and complex calibration procedures that preclude its use in consumer electronic devices or clinical settings. Several recent studies (Hwangbo et al., 2013; Wu and Pei, 2017; Olivares et al., 2009; Ren et al., 2015; Lu et al., 2022; Xia et al., 2022) proposed several calibration methods that do not need precision equipment. Specifically, in (Hwangbo et al., 2013), a camera-aided calibration method was reported. The images provide the orientation and position information of the sensor to confirm its orientation, resulting in a high computational complexity. In a separate method (Wu and Pei, 2017), a homogeneous magnetic field was employed as the calibration reference. The natural geomagnetic field is very weak and is easily affected by the alternating electric field, again making it difficult to implement outside the laboratory. In (Olivares et al., 2009; Ren et al., 2015), an accelerometer-aided gyroscope calibration method was presented. The accelerometer was first calibrated using the multi-position method. Then, the rotation speed of the sensor body is provided by the accelerometer. In this method, an extra triaxial accelerometer with the same coordinate system is needed, and the calibration error of the accelerometer affects the calibration accuracy of the gyroscope. Besides, the calibration process takes more than ten minutes. In (Lu et al., 2022), a two-step method was proposed to estimate all parameters for a six-axis IMU. A turntable was employed to perform a 12-points rotation, and a Kalman filter was designed to compensate for the parameter error. However, from the above discussion, we can infer that these gyroscope calibration methods relying on external equipment are unsuitable for scenarios where the external calibration device is unavailable, such as in a field or clinical setting akin to a busy and chaotic hospital birthing room. Besides, an accelerometer may not necessarily be integrated into the device and cannot provide calibration reference for a gyroscope. This paper, for the first time, introduced an easy-to-get rotation angle that can be used as the calibration reference. In this case, the gyroscope calibration no longer requires a constant rotation speed or reference from an accelerometer, thereby getting rid of the shackles of high-precision external equipment.
@@ 2 Methodology
In this section, we first propose an autocalibration method that uses manual rotation angle as a calibration reference. Then, a four-observation calibration method is reported. The overview of the proposed celebration method is shown in Figure. 3. In this study, we ignored the earth’s rotation as it is often submerged in the noise of the low-cost gyroscope.
@@ 2.2 Calibration principle
where the $k_{x},k_{y},k_{z}$ and $b_{x},b_{y},b_{z}$ represents the scale factors and biases, respectfully. In this paper, nonlinearity is disregarded because modern MEMS IMUs typically have nonlinearity on the order of thousandths (Li et al., 2024; TDK InvenSense, 2024), which is negligible compared to the scale factor and bias errors considered.
@@
$R_{t}=(\cos\theta_{t})I+(\sin\theta_{t})[U_{t}]_{x}+(1-\cos\theta_{t})U_{t}U_{t}^{T},$ (4)
where $I$ is a 3-by-3 identity matrix, and $[U_{t}]_{x}$ is the cross product matrix. We consider the orientation of the gyroscope to start reading (i.e. $t=0$) as the initial frame (n), and the gyroscope as body frame (b). Then, the rotation matrix $C^{n}_{b,t}$ which transforms the body frame to initial frame can be calculated as follows:
@@
where $K=\mathrm{diag}[k_{x},k_{y},k_{z}]$ contains the scale factors. Let
@@
where $\theta_{total}$ is a known value, $\tilde{C}_{t}^{n}$ is a constant value calculated by (8), and $t$ is the time of collecting $N$ samples, and $\theta_{total}$ is the total rotation angle performed in each experiment. For an embedded system, time $t$ can be measured precisely. Turning a precise angle in the experiment becomes the key to the success of the experiment.
@@
Gyroscope readings $M_{t}=[m_{x,t},m_{y,t},m_{z,t}],i\in[0,T]$ for the stationary stage ($S_{0}$) and three rotating stages ($S_{1},S_{2},S_{3}$) (See Figure. 5). $i=1$, $K^{0}=[1,1,1].$
@@ 3 Simulation
Based on the above assumptions, we formulate the following simulation process. 30 sets of parameters are randomly generated based on the assumptions above. For each set of parameters, 500 simulations were repeated. During each simulation, we created a 4-observation measurement according to the experiment scheme described in Section 2.3. For the purpose of showing that the proposed method is not sensitive to the speed variation, the rotation speed in the rotating stage was represented by a randomly generated Bézier curve (Han et al., 2008). A typical run of one simulation is shown in Figure. 6. We also generated a testing set for each simulation. The gyroscope were assumed to be rotating constantly and have equal speed projection on each axis during testing process.
@@ 4.2 Experimental Setting
calibration method turntable method (Chatfield, 1997) -aided method (Ren et al., 2015) -aided method (Wang et al., 2021)
@@
calibration method turntable method (Chatfield, 1997) -aided method (Ren et al., 2015) -aided method (Wang et al., 2021)
@@ 4.3 Comparing with existing calibration methods
To show the efficiency and accuracy of our method, we compared our method with two state-of-the-art autocalibration method (Ren et al., 2015; Wang et al., 2021) and the gold standard turntable method (Chatfield, 1997).
@@
The accelerometer-aided method. This method uses the feature that most IMUs include accelerometers, and uses accelerometers to provide a calibration reference for the calibration of the gyroscope. Following the experiment process described in (Ren et al., 2015), we reproduced the experiment using our gyroscopes. In our experiment, the inclination angle between the device and the desktop is set to 0°, 30°, and 45 °. Different from the original experiment, we did not repeat the experiment multiple times and take average to improve the calibration accuracy, because the experiment time is already as long as ten minutes.
The servomotor-aided method. This method uses a servomotor to provide a reference for calibration and provides a fast and low-cost calibration method. Due to the use of external calibration equipment, this method is not a pure autocalibration method. We have calibrated our gyroscope with reference to the calibration method provided in (Wang et al., 2021).